# Generalized Euler Equation from Effective Action: Implications for the Smarr Formula in AdS Black Holes

### v2

@@ 2 Smarr formula for AdS black holes
In thermodynamics, the Euler equation arises in the thermodynamic limit, where the volume $V$ and the number of particles $N$ both tend to infinity while maintaining a fixed ratio $\frac{N}{V}$. In this limit, the grand canonical potential becomes $\Omega=-pV$, and the Euler equation follows as $-pV=U-TS-\mu Q$. We seek to extend this reasoning to the Smarr formula. However, in the context of black hole thermodynamics, defining the thermodynamic volume $V$ and pressure $p$ is far from straightforward. This is where fluid/gravity duality offers crucial insight. To define the pressure $p$ for AdS black holes, we observe that the boundary stress-energy tensor associated with these black holes takes the form of a perfect fluid stress tensor Clifford0 ; Clifford . The boundary stress tensor for any given gravitational solution can be calculated using the formula provided in Kraus ; Clifford1 ; Skenderis :
$T_{ab}=\frac{1}{8\pi G_{N}}\left(h_{ab}\left(K-\frac{(d-2)}{l}\right)+K_{ab}+\frac{lG_{ab}}{(d-3)}+...\right)~{},$ (12)
where $h_{ab}$ is the induced metric, $K_{ab}$ is the extrinsic curvature, and $G_{ab}$ is the Einstein tensor for a fixed value of the radial direction $r$. The counterterms explicitly written in (12), are valid for $d<6$; for higher dimensions, see Clifford1 . As a reminder, $d$ is the bulk dimension, so $(d-1)$ is the boundary dimension. In our case, the boundary metric is
@@
Thus, the volume of the spatial section of the boundary spacetime is $V=l^{d-2}A_{d-2}$ where the AdS radius $l$ is acting as the radius of the $(d-2)$-dimensional sphere. For even-dimensional boundary spacetime, the conformal anomaly and Casimir energy appear Kraus ; Skenderis . Since these effects are not relevant to the present discussion, we restrict our analysis to odd-dimensional boundary spacetime. For $(d-1)=2n+1$ dimensions, the components of the boundary stress tensor are
$-T^{t}_{\;t}=\frac{(d-2)\tilde{m}}{16\pi G_{n}l^{d-2}}=\frac{M}{l^{d-2}A_{d-2}}=\frac{M}{V}~{},$ (14)
@@
Returning to the Smarr formula, substituting the conformal equation of state $M=(d-2)pV$ into (11) provides a more suggesting expression:
@@ 3.1 Review standard effective action for perfect fluids
Given the dynamical variables and symmetries of the effective action, we should, in principle, include all gauge-invariant terms in the theory. However, since our focus is on perfect fluids, we will disregard higher-derivative terms in the effective action. Moreover, as stated in the introduction, we assume that the bulk spacetime has d dimensions. Consequently, in this section, we consider a perfect fluid in an odd $(d-1)=2n+1$ dimensional manifold, ensuring that anomalies and Casimir energy do not play a role in our discussion. Finally, we limit this section to field-theoretical aspects, deferring the connection between these results and gravity to the next section.
@@
The low-energy dynamical variables for the effective action are $(d-2)$ spatial scalar fields $\sigma^{I}$ that describe the fluid element, plus a temporal scalar field $\sigma^{0}$ that acts as an internal clock for the fluid element:
@@
$\sigma^{I}\to\sigma^{I}+f^{I}(\sigma^{J})~{}.$ (22)
These symmetry transformations can be motivated by the observation that $\sigma^{I}$ represents spatial coordinates describing the fluids. As such, $\sigma^{I}$ must satisfy the comoving condition
@@
$s=\sqrt{\text{Det}(B^{IJ}G_{JK})}~{},\quad B^{IJ}=g^{ab}e_{a}^{I}e_{b}^{J}~{},$ (29)
@@
Another scalar is the chemical potential $\mu$ associated with the $U(1)$ symmetries Dubovsky2 which reads
@@
We note that both $s$ and $\mu$ are local functions of the physical spacetime where fluids is on, and they take constant values only at global thermal equilibrium.
@@
Thus, the Euler equation is an unavoidable consequence of the effective action (32). Note that the equation of state for the fluid is not used at all in this derivation.
@@
With this, we conclude our review of the effective action formulation for perfect fluids. In the following section, we will use this framework to derive a generalized Euler equation.
@@ 3.2 Generalized Euler equation
To generalize the Euler equation, we introduce a new variable into the effective action for perfect fluids. This variable must be a diffeomorphism-invariant scalar with respect to $g_{ab}$ and must remain gauge-invariant under the symmetry transformations defined in (21), (22). Furthermore, as discussed at the end of Section 2, there are scenarios in which the standard scaling argument to derive the Euler equation fails, particularly when the volume of the system is compact. In such cases, the compact space’s radius introduces an additional geometric scale to the problem. This observation suggests that the new variable in the effective action should have a geometric origin.
The gauge-invariant quantity that satisfies these criteria is
@@
The definition of this variable closely resembles the definition of the chemical potential $\mu$ presented in (31). However, the temporal scalar field $\sigma^{0}$, which represents the clock for the fluid element, differs in nature from the Stueckelberg field $\varphi$ associated with U(1) symmetry [^3] The variable $y$ has been previously studied in the literature, but Ballesteros did not clearly distinguish between the temporal scalar field $\sigma^{0}$ and the gauge scalar field $\varphi$.. Given the definition of $y$, its physical meaning may not be immediately clear. Therefore, let us first derive the generalized Euler equation and subsequently discuss the physical meaning of $y$.
@@
$\epsilon+p=sT+\mu q+y\partial_{y}\left(-\chi+s\frac{\partial\chi}{\partial y}\right)~{}.$ (46)
@@
This expression is the generalized Euler equation that we were seeking. The additional correction involves a derivative of the pressure $p$ with respect to $y$. Therefore, if the thermodynamics of a fluid in a given spacetime does not respond to variations in $y$, the additional correction $\frac{\partial p}{\partial y}$ vanishes recovering the standard Euler identity.
@@
This conservation law arises from the fact that the variable $y$, defined in (39), is invariant under spatial diffeomorphisms of the target space, $\sigma^{I}\to\sigma^{I}+f^{I}(\sigma^{J})$. Mathematically, the pair $(y,\partial_{y}p)$ is analogous to the chemical potential $\mu$ and the charge density $q$ associated with the $U(1)$ symmetry. However, it is important to emphasize that the variable $y$ was introduced to capture a geometric scale imposed by the underlying spacetime on the thermodynamics of the fluid. For this reason, unlike conventional thermodynamic variables, $y$ is not expected to admit a microscopic interpretation.
To further explore the meaning of the variable $y$, let us consider a scenario of global thermal equilibrium, where the fluid clock can be aligned with physical time; that is, $\sigma^{0}=x^{0}$. In this case, we find that $y=u^{0}$, a result that may appear somewhat opaque at first glance. To build intuition, let us perform a Weyl transformation of the physical metric, $g^{\prime}=Wg$, where $W$ is a conformal factor. Under this transformation, the fluid velocity transforms as $u^{\prime a}=\frac{u^{a}}{W}$. This leads to the following relation:
$y=u^{\prime 0}=\frac{u^{0}}{W}~{}.$ (49)
This expression suggests that the variable $y$ is naturally associated with the inverse of the conformal factor $W$. Alternatively, rather than performing an explicit Weyl transformation, one can express the physical metric as a conformal rescaling of a reference metric, $g=W\bar{g}$, leading to the same conclusion that $y\sim\frac{1}{W}$.
@@
The relation in (50) bears a notable resemblance to earlier results in the literature on effective actions for fluids Minwalla ; Jensen0  [^4] We thank Kristan Jensen for bringing these references to our attention.. However, a key distinction lies in the construction of our effective action, which is formulated at zeroth order in the derivative expansion. In contrast, the works cited in Minwalla ; Jensen0 incorporate higher-order derivative corrections, which modify the form of the free energy accordingly.
@@ 4 Smarr formula as a particular case of the generalized Euler equation
In this section, we extend the discussion from section 2 by examining the AdSd Reissner-Nordström black hole with arbitrary horizon topology. We adopt Eddington-Finkelstein coordinates, as this gauge simplifies the analysis of Weyl transformations compared to the Fefferman-Graham gauge. Moreover, we restrict the boundary dimension $(d-1)$ to be odd; otherwise, the conformal anomaly and the Casimir energy would need to be accounted for in the effective action. The line element is:
@@
Here, the gauge field $A$ is still given by equation (4). The constant $k$ represents the unit-normalized curvature of the transverse space $\Sigma_{k}$. Thus, the black brane is recovered for $k=0$, the spherical black hole for $k=1$, and the hyperbolic black hole for $k=-1$. In all cases, the black hole energy $M$ and charge $Q$ are related to the parameters $\tilde{m}$ and $\tilde{q}$ as follows:
@@
For for $k=0$ the last term vanishes, thereby recovering the standard Euler equation for a charged black brane. This is because, in flat space, there are no additional scales to consider.
We will now compute the thermodynamic potential $\chi$ (40) which is associated with the effective action for the holographic perfect fluid dual to the AdSd Reissner-Nordström black hole. Our objective is to place the holographic fluid in a boundary spacetime that exhibits spacetime dependence, allowing us to mimic a state that is near out of equilibrium [^5] This step is purely for illustrative purposes and carries no additional meaning, as we are disregarding dissipation, and anomalies.. As a consequence, the thermodynamic variables now will become local functions of the spacetime. To obtain the boundary metric with the desired properties, we perform the following radial diffeomorphism in the bulk line element (51):
@@
This line element describes a non-homogeneous, time-dependent background in which the fluid resides. The radial diffeomorphism (56) consequently induces a Weyl transformation on the boundary metric [^6] In other words, the Weyl transformation is induced by a PBH transformation of the bulk metric. A detailed discussion of the consequences of this transformation in the thermodynamics of asymptotically locally AdS black holes is given in Ioannis3 ., leading to the following expression for the fluid velocity:
@@
that matches the horizon metric  [^7] More precisely, the horizon metric before the Weyl transformation.. Given the target metric and the fluid velocity, we can use the formulas (29), (31), and (39) to compute the entropy density $s$, the chemical potential $\mu$, and the variable $y$, respectively:
$s=\frac{\Sigma_{k,d-2}}{4G_{N}}\frac{r_{h}^{d-2}}{W^{d-2}}~{},\quad\mu=\frac{\tilde{q}}{\gamma r_{h}^{d-3}}\frac{1}{W}~{},\quad y=\frac{l}{W}~{},$ (62)
which are the variables that the thermodynamic potential $\chi$ depends on. We note that the entropy density satisfies its conservation law (37) as expected. We will utilize these variables to derive the explicit expression for the new thermodynamic potential $\chi$ from the energy density $\epsilon$ using equation (42). Our next step involves computing the boundary stress tensor via the formula (12) and the boundary metric given in (59). The components of the stress tensor are
$-T^{t}_{\;t}=\frac{lM}{W^{d-1}\Sigma_{k,d-2}}~{},$ (63)
@@
$\epsilon=\frac{(d-2)}{16\pi G_{N}l}\left[\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-3}{d-2}}(ky^{2}+(\gamma l)^{2}\mu^{2})+\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-1}{d-2}}\right]~{},$ (65)
@@
$\chi=\frac{(d-2)}{16\pi G_{N}l}\left[-\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-3}{d-2}}(ky^{2}+(\gamma l)^{2}\mu^{2})+\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-1}{d-2}}\right]~{}.$ (67)
@@
$T=\frac{(d-1)}{4\pi l}\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{1}{d-2}}+\frac{(d-3)(ky^{2}-(\gamma l)^{2}\mu^{2})}{4\pi l}\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{-1}{d-2}}~{},$ (68)
$q=\frac{(d-2)\gamma^{2}l\mu}{8\pi G_{N}}\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-3}{d-2}}~{}.$ (69)
@@
$\frac{\partial p}{\partial y}=\frac{ky}{8\pi G_{N}l}\left(\frac{4G_{N}s}{\Sigma_{k,d-2}}\right)^{\frac{d-3}{d-2}}~{}.$ (70)
@@
Furthermore, substituting the expressions for $y$ and $\mu$ from (62), one can confirm that $q$ [^8] Let us emphasize that $\tilde{q}$ denotes the constant appearing in the black hole line element, whereas $q=q(x)$ is a local function defined on the boundary spacetime. We apologize for any notational inconvenience this may cause. and $\partial_{y}p$ satisfy their respective conservation laws, given by (38) and (48).

### v3

@@ 2 Smarr formula for AdS black holes
In thermodynamics, the Euler equation arises in the thermodynamic limit, where the volume $V$ and the number of particles $N$ both tend to infinity while maintaining a fixed ratio $\frac{N}{V}$. In this limit, the grand canonical potential becomes $\Omega=-pV$, and the Euler equation follows as $pV=U-TS-\mu Q$. We seek to extend this reasoning to the Smarr formula. However, in the context of black hole thermodynamics, defining the thermodynamic volume $V$ and pressure $p$ is far from straightforward. This is where fluid/gravity duality offers crucial insight. To define the pressure $p$ for AdS black holes, we observe that the boundary stress-energy tensor associated with these black holes takes the form of a perfect fluid stress tensor Clifford0 ; Clifford . The boundary stress tensor for any given gravitational solution can be calculated using the formula provided in Kraus ; Clifford1 ; Skenderis :
[…]
where $h_{ab}$ is the induced metric, $K_{ab}$ is the extrinsic curvature, and $G_{ab}$ is the Einstein tensor for a fixed value of the radial direction $r$. The counterterms explicitly written in (12), are valid for $d<6$; for higher dimensions, see Clifford1 . As a reminder, $d$ is the bulk dimension, so (d1)(d-1) is the boundary dimension. In our case, the boundary metric is
@@
Thus, the volume of the spatial section of the boundary spacetime is $V=l^{d-2}A_{d-2}$ where the AdS radius $l$ is acting as the radius of the (d2)(d-2)-dimensional sphere. For even-dimensional boundary spacetime, the conformal anomaly and Casimir energy appear Kraus ; Skenderis . Since these effects are not relevant to the present discussion, we restrict our analysis to odd-dimensional boundary spacetime. For (d1)=2n+1(d-1)=2n+1 dimensions, the components of the boundary stress tensor are
$T^{t}_{\;t}=\frac{(d-2)\tilde{m}}{16\pi G_{n}l^{d-2}}=\frac{M}{l^{d-2}A_{d-2}}=\frac{M}{V}~,$ (14)
@@
Returning to the Smarr formula, substituting the conformal equation of state M=(d2)pVM=(d-2)pV into (11) provides a more suggesting expression:
@@ 3.1 Review standard effective action for perfect fluids
Given the dynamical variables and symmetries of the effective action, one should, in principle, include all gauge-invariant terms organized in a derivative expansion. Since our focus is restricted to perfect fluids, however, we will neglect the higher-derivative contributions that encode dissipative effects. Moreover, as stated in the introduction, we assume that the bulk spacetime has d dimensions. Consequently, in this section, we consider a perfect fluid in an odd (d1)=2n+1(d-1)=2n+1 dimensional manifold, ensuring that anomalies and Casimir energy do not play a role in our discussion. Finally, we limit this section to field-theoretical aspects, deferring the connection between these results and gravity to the next section.
@@
The low-energy dynamical variables for the effective action are (d2)(d-2) spatial scalar fields $\sigma^{I}$ that describe the fluid element, plus a temporal scalar field $\sigma^{0}$ that acts as an internal clock for the fluid element:
@@
$\sigma^{I}\to f^{I}(\sigma^{J})~.$ (22)
These symmetry transformations can be motivated by the observation that $\sigma^{I}$ represents spatial coordinates describing the fluids. As such, $\sigma^{I}$ must satisfy the comoving condition [^2] In the literature, this condition defined the Lagrangian observer.
@@
$s=\tau_{0}\sqrt{\text{Det}(B^{IJ}G_{JK})}~,\quad B^{IJ}=g^{ab}e_{a}^{I}e_{b}^{J}~,$ (29)
@@
Here, $\tau_{0}$ is a dimensionful constant that provides the correct unit for the entropy density. Another scalar is the chemical potential $\mu$ associated with the $U(1)$ symmetries Dubovsky2 which reads
@@
We note that both $s$ and $\mu$ are local functions of the physical spacetime where fluid is on, and they take constant values only at global thermal equilibrium.
@@
Thus, the Euler equation is an unavoidable consequence of the effective action (32). Note that no equation of state for the fluid is used at all in this derivation.
@@
Finally, the entropy current can be rewritten using the Euler equation (36) as
$s^{a}=p\beta^{a}-T^{ab}\beta_{b}-\frac{\mu}{T}q^{a}~,$ (39)
where $\beta^{a}\equiv\frac{u^{a}}{T}$. This expression for $s^{a}$ can serve as a starting point for the current formalism of fluids Ranga5 . With this, we conclude our review of the effective action formulation for perfect fluids. In the following section, we will use this framework to derive a generalized Euler equation.
@@ 3.2 Generalized Euler equation
To generalize the Euler equation, we introduce a new variable into the effective action for perfect fluids. This variable must be a diffeomorphism-invariant scalar with respect to $g_{ab}$ and must remain gauge-invariant under the shift symmetries given in (21), (22). As discussed at the end of Section 2, there are scenarios in which the standard scaling argument used to derive the Euler equation fails—for instance, when the volume of the system is compact. In such a case, the radius of the compact space introduces an additional geometric scale into the problem, implying that different thermodynamic quantities respond differently under scale transformations. This observation suggests that the new variable we seek should have a geometric origin and transform appropriately under scaling. The gauge-invariant quantity that satisfies these criteria is
@@
The definition of this variable closely resembles the definition of the chemical potential $\mu$ presented in (31). However, the temporal scalar field $\sigma^{0}$, which represents the clock for the fluid element, differs in nature from the Stueckelberg field $\varphi$ associated with U(1) symmetry. [^4] The variable $y$ has been previously considered in the literature, but Ballesteros did not clearly distinguish between the temporal scalar field $\sigma^{0}$ and the gauge scalar field $\varphi$. In fact, another way to understand the expression in (40) is to notice that the pushforward of the fluid velocity is given by the vector
$Y^{M}=u^{a}\partial_{a}\sigma^{M}$ (41)
which lives in the fluid space. However, due to the shift symmetries (21) and (22), only the temporal component of $Y^{M}$ is non-vanishing. Hence, the expression in (40) simply reduces to $y\equiv Y^{0}$. Let us note that the vector $Y^{M}$ of the fluid space transforms covariantly under arbitrary diffeomorphisms of the target space. Under the restricted shift symmetries, however, its temporal component remains gauge-invariant. Since the target space is a null surface, this makes the variable $y$ a natural quantity to introduce within the fluid EFT. These observations clarify the mathematical origin of $y$; its physical interpretation, on the other hand, is less immediate. We will return to this point at the end of the section, where the physical meaning of $y$ will be analyzed in detail.
@@
$\epsilon+p=sT+\mu q+y\partial_{y}\left(-\chi+s\frac{\partial\chi}{\partial s}\right)~.$ (48)
@@
This expression is the generalized Euler equation that we were seeking. The additional correction involves a derivative of the pressure $p$ with respect to $y$. Therefore, if the thermodynamics of a fluid in a given spacetime does not respond to variations in $y$, the additional correction $\frac{\partial p}{\partial y}$ vanishes, recovering the standard Euler identity.
@@
This conservation law arises from the fact that the variable $y$, defined in (40), is invariant under spatial diffeomorphisms of the target space, $\sigma^{I}\to f^{I}(\sigma^{J})$. Because of this, the entropy current (39) is now modified to be
$s^{a}=p\beta^{a}-T^{ab}\beta_{b}-\frac{\mu}{T}q^{a}-\frac{y}{T}J^{a}_{(y)}~.$ (51)
To further explore the physical meaning of the variable $y$, let us consider a scenario of global thermal equilibrium, where the fluid clock can be aligned with physical time; that is, $\sigma^{0}=x^{0}$. In this case, we find that $y=u^{0}$, an expression that may appear somewhat opaque from the physics point of view at first glance. To build intuition, let us perform a Weyl transformation of the physical metric, $g^{\prime}=W^{2}g$, where $W$ is a conformal factor. Under this transformation, the fluid velocity transforms as $u^{\prime a}=\frac{u^{a}}{W}$. This leads to the following relation:
$y\to y^{\prime}=\frac{u^{0}}{W}~.$ (52)
Thus, the variable $y$ is naturally associated with the inverse of the conformal factor $W$. To fully grasp the physical meaning of the variable $y$, let us recall that thermodynamic variables are not Weyl invariant; the introduction of the variable $y$ provides a convenient bookkeeping device to track how these variables transform under scaling transformations. In fact, for constant $y$ the operator $y\partial_{y}$ appearing in the generalized Euler equation (49) plays the role of the dilation operator, since $y\partial_{y}=x^{a}\partial_{a}$ in line with the interpretation discussed above. Certainly, for non-constant $y$, the operator $y\partial_{y}$ should not be identified anymore with the dilation operator.
@@
The relation in (53) bears a notable resemblance to earlier results in the literature on effective actions for fluids Minwalla ; Jensen0 . [^5] We thank Kristan Jensen for bringing these references to our attention. However, a key distinction lies in the construction of our effective action, which is formulated at zeroth order in the derivative expansion. In contrast, the works cited in Minwalla ; Jensen0 incorporate higher-order derivative corrections, which modify the form of the free energy accordingly.
@@ 4 Smarr formula as a particular case of the generalized Euler equation
In this section, we extend the discussion from section 2 by examining the AdSd Reissner-Nordström black hole with arbitrary horizon topology. We adopt Eddington-Finkelstein coordinates, as this gauge simplifies the analysis of Weyl transformations compared to the Fefferman-Graham gauge. Moreover, we restrict the boundary dimension (d1)(d-1) to be odd; otherwise, the conformal anomaly and the Casimir energy would need to be accounted for in the effective action. The line element is:
@@
Here, the gauge field $A$ is still given by equation (4). The constant $k$ represents the unit-normalized curvature of the transverse space $\Sigma_{k}$. Thus, the black brane is recovered for $k=0$, the spherical black hole for $k=1$, and the hyperbolic black hole for k=1k=-1. In all cases, the black hole energy $M$ and charge $Q$ are related to the parameters $\tilde{m}$ and $\tilde{q}$ as follows:
@@
For $k=0$ the last term vanishes, thereby recovering the standard Euler equation for a charged black brane. This is because, in flat space, there are no additional scales to consider.
We will now compute the thermodynamic potential $\chi$ (42) which is associated with the effective action for the holographic perfect fluid dual to the AdSd Reissner-Nordström black hole. Our objective is to place the holographic fluid in a boundary spacetime that exhibits spacetime dependence, allowing us to mimic a state that is near out of equilibrium. [^6] This step is purely for illustrative purposes and carries no additional meaning, as we are disregarding dissipation, and anomalies. As a consequence, the thermodynamic variables now will become local functions of the spacetime. To obtain the boundary metric with the desired properties, we perform the following radial diffeomorphism in the bulk line element (54):
@@
This line element describes a non-homogeneous, time-dependent background in which the fluid resides. The radial diffeomorphism (59) consequently induces a Weyl transformation on the boundary metric, [^7] In other words, the Weyl transformation is induced by a PBH transformation of the bulk metric. A detailed discussion of the consequences of this transformation in the thermodynamics of asymptotically locally AdS black holes is given in Ioannis3 . leading to the following expression for the fluid velocity:
@@
that matches the horizon metric. [^8] More precisely, the horizon metric before the Weyl transformation. Given the target metric and the fluid velocity, we can use the formulas (29), (31), and (40) to compute the entropy density $s$, the chemical potential $\mu$, and the variable $y$, respectively:
$s=\frac{1}{4G_{N}}\frac{r_{h}^{d-2}}{W^{d-2}}~,\quad\mu=\frac{\tilde{q}}{\gamma r_{h}^{d-3}}\frac{1}{W}~,\quad y=\frac{l}{W}~,$ (65)
which are the variables that the thermodynamic potential $\chi$ depends on. We note that the entropy density satisfies its conservation law (37), as expected, and we set $\tau_{0}=\frac{1}{4G_{N}}$ in equation (29). We will utilize these variables to derive the explicit expression for the new thermodynamic potential $\chi$ from the energy density $\epsilon$ using equation (44). Our next step involves computing the boundary stress tensor via the formula (12) and the boundary metric given in (62). The components of the stress tensor are
$T^{t}_{\;t}=\frac{lM}{W^{d-1}\Sigma_{k,d-2}}~,$ (66)
@@
$\epsilon=\frac{(d-2)}{16\pi G_{N}l}\left[\left(4G_{N}s\right)^{\frac{d-3}{d-2}}(ky^{2}+(\gamma l)^{2}\mu^{2})+\left(4G_{N}s\right)^{\frac{d-1}{d-2}}\right]~,$ (68)
@@
$\chi=\frac{(d-2)}{16\pi G_{N}l}\left[-\left(4G_{N}s\right)^{\frac{d-3}{d-2}}(ky^{2}+(\gamma l)^{2}\mu^{2})+\left(4G_{N}s\right)^{\frac{d-1}{d-2}}\right]~.$ (70)
@@
$T=\frac{(d-1)}{4\pi l}\left(4G_{N}s\right)^{\frac{1}{d-2}}+\frac{(d-3)(ky^{2}-(\gamma l)^{2}\mu^{2})}{4\pi l}\left(4G_{N}s\right)^{\frac{-1}{d-2}}~,$ (71)
$q=\frac{(d-2)\gamma^{2}l\mu}{8\pi G_{N}}\left(4G_{N}s\right)^{\frac{d-3}{d-2}}~.$ (72)
@@
$\frac{\partial p}{\partial y}=\frac{ky}{8\pi G_{N}l}\left(4G_{N}s\right)^{\frac{d-3}{d-2}}~.$ (73)
@@
Furthermore, substituting the expressions for $y$ and $\mu$ from (65), one can confirm that qq [^9] Let us emphasize that $\tilde{q}$ denotes the constant appearing in the black hole line element, whereas $q=q(x)$ is a local function defined on the boundary spacetime. We apologize for any notational inconvenience this may cause. and $\partial_{y}p$ satisfy their respective conservation laws, given by (38) and (50).